y^2/3 + 2y^1/3 -3 = 0
Solve For Y

Answer by htmentor(1343)

Let x = y^1/3
Then we can write the equation as
x^2 + 2x - 3 = 0
This can be factored as
(x-1)(x+3) = 0
The solutions are x=1, x=-3
So we have y^1/3 = 1 and y^1/3 = -3
y^1/3 = 1 -> y = (1)^3 = 1
y^1/3 = -3 -> y = (-3)^3 = -27
Ans: y=1, y=-27